Direct Labor and Direct Materials Variances, Journal Entries

Jameson Company produces paper towels. The company has established the following direct materials and direct labor standards for one case of paper towels:

Paper pulp (3 lbs. @ $0.40) $ 1.20
Labor (2 hrs. @ $12) 24.00
    Total prime cost $25.20

During the first quarter of the year, Jameson produced 55,000 cases of paper towels. The company purchased and used 165,700 pounds of paper pulp at $0.37 per pound. Actual direct labor used was 111,000 hours at $12.10 per hour.

Required:

Instructions for parts 1 and 2: If a variance is zero, enter "0" and select "Not applicable" from the drop down box.

1. Calculate the direct materials price and usage variances.

Materials Price Variance $
Materials Usage Variance $

2. Calculate the direct labor rate and efficiency variances.

Labor Rate Variance $
Labor Efficiency Variance $

3. Prepare the journal entries for (1) the direct materials price variance, (2) the direct materials usage variance, and (3) the direct labor variances. If an amount box does not require an entry, leave it blank or enter "0".
Answer #1

Solution:

1. Calculation of direct material price and usage variance.

Material price variance = (Standard price - Actual price ) x Actual quantity

= (0.4 - 0.37) x 165,700

= 4,971 Favorable

Material usage variance = (Standard quantity - Actual Quantity) x Standard price

= [ (55,000 units x 3 lbs per unit) - 165,700] x 0.4

= ( 165,000 - 165,700) x 0.4

= 280 Unfavorable

2. Calculation of labor rate and efficiency variance

Labor rate variance = (Standard rate - actual rate) x actual hours worked

= (12 - 12.1) x 111,000

= 11,100 Unfavorable

Labor efficiency variance = (Standard hours - Actual Hours ) x Standard rate

= [( 55,000 units x 2 hrs) - 111,000] x 12

= 12,000 Unfavorable

3.

Sl.no Journal Debit Credit
1. Materials ( 165,700 lbs x 0.4) 66,280
Material Price Variance ( 165,700 x 0.03) 4,971
Accounts Payable ( 165,700 lbs x 0.37) 61,309
( Being material purchased recorded at standard cost and balance between standard and actual price recorded as material price variance)
2. Work in process ( 165,000 x 0.4 ) 66,000
Materials Usage Variance ( 700 x 0.4) 280
Materials ( 165,700 x 0.4) 66,280
( Being materials issued for production in standard quantity and balance quantity recorded as material usage variance)
3. Work in process (110,000 hrs x 12 ) 1,320,000
Labor rate variance ( 111,000 hrs x 0.1) 11,100
Labor efficiency variance ( 1000 hrs x 12) 12,000
Wages payable ( 111,000 hrs x 12.1) 1,343,100
(Being wages payable recorded at actual rate and labor used for production recorded at standard rates and balance recorded as variances as calculated above)